# From Frege to chatGPT: Compositionality in language, cognition, and deep neural networks

### v1

@@ 4.1 Operationalizing Compositionality
However, recent work has investigated methods that endow neural networks with inductive biases that would facilitate their adaptation to these shifts and allow them to replicate the behavioral signatures of compositionality. In the following, we’ll review recent work from three different areas suggesting that neural networks can acquire such inductive biases:
  1. 1.
    Architectural inductive biases: Some work has focused on designing specialized network architectures that would facilitate the emergence of compositional generalization behavior. This can include neuro-symbolic hybrids, which combine explicit symbolic processes with neural networks, or specialized neural architectures that do not build in such explicit symbolic processes.
  2. 2.
    Metalearning: In the metalearning setting (Lake and Baroni,, 2023), an inductive bias is imparted to a neural network by training it on a specific distribution of tasks (e.g., compositional generalization tasks), endowing it with an ability to learn new tasks in specific ways.
  3. 3.
    Large-scale pretraining: When neural networks are pretrained on a large amount of data, they can acquire emergent capabilities (Brown et al.,, 2020; Bubeck et al.,, 2023). Similar to the metalearning setting, pretraining can be thought of as a way of imparting inductive biases to a learner.
@@
In the following, we will review key findings from recent studies exploring compositionality using each of these approaches, highlighting how each challenges the traditional view of compositionality in neural networks. Our review will largely focus on language, as this has been the domain in which compositionality has traditionally played a central explanatory role (see section 2.1). However, there has also been a recent surge of interest in compositionality in vision (Lepori et al.,, 2023; Zhou et al.,, 2024), multi-modal models (Castro et al.,, 2024; Hsieh et al.,, 2023; Lewis et al.,, 2023; Ma et al., 2023b, ; Ma et al., 2023a, ; Mitra et al.,, 2023; Ossowski et al.,, 2024; Shukor et al.,, 2023), and in reinforcement learning (Bakirtzis et al.,, 2022; Jothimurugan et al.,, 2021; Liu and Frank,, 2022; Žikelić et al.,, 2023), where similar issues have been explored. Furthermore, while we have chosen to focus on the above three topics to organize our review, it should be noted that much of the recent work does not fall cleanly into any of these categories. For example, some work has investigated the effects of standard architectural changes on compositional generalization (Csordás et al.,, 2022), explored techniques involving data augmentation (Andreas,, 2020; Cazzaro et al.,, 2024; Chai et al.,, 2023; Jiang et al.,, 2022), injecting structural knowledge via additional supervision or auxiliary tasks (Chakravarthy et al.,, 2022; Jiang and Bansal,, 2021), and investigated the limits of standard transformers on compositional tasks (Dziri et al.,, 2023; Zhou et al., 2023b, ).
@@
After the review, we discuss how these new approaches may inform our understanding of human compositionality, including their implications for its neural mechanisms and development. We conclude with a discussion of the ‘mere implementation’ objection, the second horn of Fodor and Pylyshyn’s (1988) argument.
4.2 Architectural Inductive Biases
A major focus of the work on compositionality in neural networks has been the search for architectural inductive biases that would allow neural networks to generalize on compositional tests like the SCAN task described above. The inductive biases that have been explored vary in the extent to which they explicitly incorporate the exact kind of compositionality characteristic of classical or symbolic systems. Neuro-symbolic hybrid architectures incorporate explicit symbolic operations into neural network models, while strictly neural architectures attempt to implement inductive biases that could plausibly facilitate the emergence of structure-sensitive, compositional processing without the use of such explicit symbolic operations. [^19] The term “neuro-symbolic” is sometimes used to describe both of these kinds of approaches together—those that attempt to integrate explicit symbolic representations or operations in a neural network architecture as well as the approaches that take inspiration from symbolic or classical computation to design strictly neural architectures. However, we believe the distinction between these approaches is important, and therefore reserve the term “neuro-symbolic hybrid” for the former. This distinction has also been usefully described as a difference between “neuro-compositional” and “neuro-symbolic” or between “neurons-in-structure” and “structure-in-neurons” (Smolensky et al., 2022b, ). Each of these approaches is a very active area of research, so a comprehensive review is beyond the scope of this chapter, but in the following we provide a targeted survey (see Yu et al.,, 2023; Smolensky et al., 2022b, ; Webb et al.,, 2024, for other reviews).
Neuro-symbolic hybrids (Ellis et al.,, 2021; Gaur et al.,, 2022; Marcus,, 2020; Marcus and Davis,, 2020; Yu et al.,, 2023; Zhou et al.,, 2024) are systems that attempt to fuse neural and symbolic modules or components into a single system that garners the benefits of each—the power and flexibility demonstrated by modern deep learning systems in dealing with high-dimensional inputs like text and images (LeCun et al.,, 2015), combined with the strength of symbolic modules in ensuring that representations or processes are compositional (Fodor and Pylyshyn,, 1988; Marcus,, 1998).
For example, Neural Module Networks (Andreas et al.,, 2017, 2016; Bahdanau et al.,, 2020) assign separate neural “modules” to specific primitive operations (e.g., filtering or querying an image based on certain attributes like color or shape, or checking if two colors are the same). These modules can be composed like symbolic functions to dynamically construct a network to accomplish a specific series of operations. For example, in the visual-question answering setting, they can be used to answer questions like: “Do the green object to the left of the red sphere and the blue object to the right of the orange cylinder have the same shape?" (Johnson et al.,, 2016). When they are tested in a compositional generalization setting, they show some advantages relative to standard networks, but still appear to be inferior to humans (Bahdanau et al.,, 2018, 2020).
In another neuro-symbolic approach, Neural Program Synthesis (Nye et al., 2020a, ; Xu et al.,, 2019; Zhong et al.,, 2023), building on work in program synthesis (Gulwani et al.,, 2017) and probabilistic program induction (Lake et al.,, 2015), explicit symbolic rules or programs are inferred from data using neural networks. These explicit rules can then be manipulated or processed using traditional programming languages to harness the compositional power of these classical systems. On the SCAN task, for example, a neuro-symbolic model based on neural program synthesis (Nye et al., 2020a, ) was able to infer the correct grammar from only a few examples and generalize to given test samples, thereby performing near-perfectly on the compositional split. In general, neuro-symbolic models have demonstrated an impressive ability to learn visual and linguistic concepts from data, and to generalize these to unseen examples (Ellis et al.,, 2021; Mao et al.,, 2019).
It is important to note that while the neuro-symbolic approach may show that it is possible for neural networks to implement compositional generalization behaviors when they are augmented with explicit symbolic processes, it cannot show that the neural components themselves offer an alternative explanation of these generalization behaviors. In other words, these systems are designed to ‘merely implement’ the same kind of compositionality present in classical symbolic systems, and so do not so much grapple with the second horn of the compositionality challenge (2.3), as embrace it (for discussion, see 5.3).
Work investigating strictly neural architectural inductive biases, on the other hand, seeks to meet this challenge by showing that compositional generalization does not require explicit rules, symbols or hand-coded representations (Altabaa et al.,, 2023; Kriete et al.,, 2013; Palangi et al.,, 2017; Rougier et al.,, 2005; Russin et al., 2020a, ; Soulos et al.,, 2023; Schlag et al.,, 2019; Webb et al.,, 2024, 2023). [^20] Again, this approach has been referred to as “neurocompositional” or “structure-in-neurons,” as opposed to the “neuro-symbolic” or “neurons-in-structure” approach (Smolensky et al., 2022b, ). The goal of this work is to find an architectural inductive bias generic enough that it can be applied across many domains (e.g., language tasks, vision tasks, etc.), while still empowering DNNs to generalize in new ways (Battaglia et al.,, 2018). If such an architectural feature is discovered, it may shed light on how the architecture of the human brain could have evolved a bias for compositionality without building in overt symbolic operations, which may have been unlikely to emerge in a biological system (O’Reilly et al.,, 2014).
Some of this work has taken direct inspiration from symbolic systems, however, and attempted to translate compositional features of classical computing, like role-filler independence (Chakravarthy et al.,, 2022; Palangi et al.,, 2017; Schlag et al.,, 2019; Smolensky,, 1990) and variable binding (Webb et al.,, 2021), into architectural features that could plausibly be encoded in a strictly neural system. Other work has tried to replicate some of the architectural motifs known to exist in the brain, like those associated with the prefrontal cortex (Russin et al., 2020b, ), the parietal cortex (O’Reilly et al.,, 2021; Summerfield et al.,, 2020), or the medial temporal lobe (Behrens et al.,, 2018; Whittington et al.,, 2020).
Figure 4: Architectural inductive biases. Modifying a standard encoder-decoder architecture (Bahdanau et al.,, 2014) can instantiate a different set of inductive biases, which can have important implications for how the model generalizes. One such bias—the separation of representations for structure and content (O’Reilly et al.,, 2021; Russin et al., 2020a, )—can facilitate compositional generalization on the SCAN task, transforming an o.o.d. generalization problem into two separate i.i.d. generalization problems (one involving a known structure, such as the relationship between adverbs and verbs, and one involving known content, such as the meaning of the primitive verb “jump”). Figure adapted from Russin et al., 2020a .
One major theme of this work has been the use of architectural motifs that encourage the model to learn relatively independent representations for structural (e.g., “role”) information and content (e.g., “filler”) information. The idea here is that if structural information (e.g., general knowledge about verbs or adverbs in a given language) is learned separately from content information (e.g., knowledge about a specific verb like “walk”), then this will allow abstract, content-independent structural knowledge to be generalized to other specific contents (e.g., “jump”).
For example, tensor product representations (Smolensky,, 1990), in which vectors representing “role” information and vectors representing “filler” information are bound by the tensor product operation to construct composite representations, has been extended to the deep learning framework in a number of different ways (Jiang et al.,, 2021; McCoy et al.,, 2019; Palangi et al.,, 2017; Schlag et al.,, 2019; Soulos et al.,, 2020, 2022; Smolensky et al., 2022a, ). The original work with tensor products (Smolensky,, 1990) sought to establish how a strictly neural system could embody the principle of compositionality and represent constituent structure (Fodor and Pylyshyn,, 1988), but was unconcerned with the problem of learning such representations from data. However, this newer work explores tensor product representations in the context of learning by building tensor-product-like operations into DNN architectures such as RNNs (Palangi et al.,, 2017) or transformers (Schlag et al.,, 2019), and measuring the extent to which it encourages the models to represent constituent structure in an interpretable, compositional way (Russin et al., 2021a, ; Soulos et al.,, 2020; Smolensky et al., 2022a, ). These networks have proven capable of performing well on tasks such as mathematical word problems (Schlag et al.,, 2019), machine translation (Soulos et al.,, 2022), and abstractive summarization (Jiang et al.,, 2021), though they have been shown to fail in some cases of o.o.d. generalization and extrapolation (e.g., in the math setting; Schlag et al.,, 2019).
Other work has implemented similar ideas about role-filler independence and variable-binding, but without the use of explicit tensor product representations (Chakravarthy et al.,, 2022; Russin et al., 2020a, ; Jiang et al.,, 2024; Webb et al.,, 2024). For example, Russin et al., 2020a showed that when a strict separation between structure-related information and content-related information is imposed in the attention mechanism of an encoder-decoder RNN architecture (Bahdanau et al.,, 2014), performance on the compositional generalization test in the SCAN task is improved to 78.4% accuracy. Concurrent work based on a similar principle showed that this performance can be improved even further to 98.8% when a few small changes, such as adding noise, are introduced (Li et al.,, 2019). The structure-content separation in these models is imposed not by hand-coding specific representations, but by modifying the architecture so that separate information-processing streams are responsible for (1) choosing what to attend in the input at each time step (we can label this the “structure” stream), and (2) for translating the individual instruction-words that are currently attended into their corresponding actions (we can label this the “content” stream). This modification forces the model to “factorize” (Behrens et al.,, 2018; Whittington et al.,, 2020) the problem into two sub-problems: the first is to find the right attentional procedure or routine (i.e., the right sequence of decisions about where to attend at each time point), and the second is to memorize how individual primitive verbs should be translated into their corresponding actions. In effect, this factorization turns an o.o.d. problem into two i.i.d problems (see Figure 4). Rather than confronting a novel structure-content combination that lies outside of its training distribution, the model can combine two in-distribution tasks: one where the individual verb is translated into its corresponding action (i.i.d. with respect to training because “jump” → JUMP appears in the training set), and one where the routine for a given grammatical structure is executed (i.i.d. with respect to the model’s training because all constructions with other primitive verbs appear in the training set).
An alternative way of thinking about this kind of separation is as a “relational bottleneck” (Altabaa et al.,, 2023; Campbell and Cohen,, 2024; Campbell et al.,, 2023; Webb et al.,, 2024): these architectural motifs force the network (or part of the network, e.g., the “role” or “structure” stream) to represent information that is purely relational, or independent of any content associated with individual words or objects. This encourages compositionality because it allows generalization to novel content-relation or content-structure combinations, as long as the individual contents and relations are familiar (i.e., they are i.i.d. with respect to the training data). These ideas have been explored in a number of different settings, including in the Emergent Symbol Binding Network capable of solving a task based on Raven’s Progressive Matrices (Webb et al.,, 2021) and in applications to abstract problem-solving tasks (Altabaa et al.,, 2023).
Both neuro-symbolic hybrids and strictly neural architectural inductive biases have proven successful in exhibiting compositional generalization behaviors in limited domains such as the SCAN dataset and related tasks (Nye et al., 2020b, ; Russin et al., 2020a, ; Ossowski et al.,, 2024). Both approaches can be understood as succeeding on o.o.d. generalization problems by instantiating powerful inductive biases in learning systems: neuro-symbolic models leverage overtly classical elements such as symbolic or rule-like representations, whereas neural architectural inductive biases attempt to realize similar properties in strictly neural systems. These inductive biases can in some cases turn what looks like an o.o.d. generalization problem into multiple separate i.i.d. generalization problems (see Figure 4; Russin et al., 2020a, ).
While both of these approaches are active areas of research, there has been some difficulty in scaling them up to difficult tasks such as language modeling, where large, unbiased transformer architectures trained on huge datasets have dominated (Bubeck et al.,, 2023; Brown et al.,, 2020). This may be because it is more difficult to design strong inductive biases (whether neuro-symbolic or strictly neural) by hand than it is to learn them from data, a familiar dynamic encapsulated in the “bitter lesson” (Sutton,, 2019). The remaining sections of this review discuss work exploring whether such data-driven methods, including metalearning and language-model pretraining, can also endow neural networks with the inductive biases required for compositional generalization.
4.3 In-Context Compositionality via Metalearning
In the metalearning paradigm (Binz et al.,, 2023; Griffiths et al.,, 2019; Sandbrink and Summerfield,, 2024; Wang et al.,, 2018; Wang,, 2020), rather than building inductive biases into a network’s architecture, they are imparted to the model by specifying a distribution of tasks on which it is trained. This has the advantage that the specific architectural features necessary to achieve the learning behaviors of interest do not need to be hand-designed in advance (McCoy et al.,, 2020). Instead, the researcher (who has learned their “bitter lesson”; Sutton,, 2019) can simply define the set of learning problems at which the model should excel and rely on the optimization process during metalearning to discover whatever inductive biases are useful for accomplishing these behaviors. Thus, metalearning neural networks learn how to learn—they are trained to rapidly acquire new skills or generalize in specific ways on new problems.
Figure 5: The inner-loop learning algorithm that emerges in a network’s activation dynamics may have different properties than the original learning algorithm used to train the network’s weights. These two learning algorithms can simultaneously coexist in a single network.
@@ 4.3 In-Context Compositionality via Metalearning
A number of metalearning methods have been developed for use with neural networks (Bengio et al.,, 1991, 1995; Duan et al.,, 2016; Finn et al.,, 2017; Hospedales et al.,, 2020; Schmidhuber,, 1987). In many of these methods, the model learns to implement in its forward activation dynamics a separate “inner-loop” or “in-context” learning (ICL) algorithm, distinct from the “outer-loop” or “in-weight” learning (IWL) algorithm used to train the network in the first place (see Figure 5). In this case, the network’s inner-loop algorithm “learns” within the flow of information from its inputs to its outputs (Duan et al.,, 2016; Lake,, 2019; Lake and Baroni,, 2023; Santoro et al.,, 2016; Wang et al.,, 2018). The network’s weights are trained such that the forward pass of the model can take some study examples as inputs in context and learn from them to produce the correct generalization behavior on a test query (see Figure 6). For example, a sequence-processing network such as an RNN or a transformer can be trained on inputs that consist of (1) a set of study examples along with their labels, and (2) a single test query whose label is not given, but must be inferred. In this case, the network is trained to produce a prediction about the label corresponding to the test query, conditioned on the set of training samples and their labels. The model is then evaluated by freezing its weights and testing on new study examples and a new test query drawn from another task.
Figure 6: Metalearning setting. (A) In a typical (non-metalearning) training scenario, a learner presented with a compositional generalization problem is trained on the training set (green, numbered 1, 2, 3) and then tested on the test set (blue). In this case, compositional generalization requires o.o.d. generalization. (B) In a metalearning scenario, the metalearner is trained on a distribution of tasks (“episodes”). In each task, the metalearner is presented with a set of in-context study examples (green) that are analogous to the training set in the vanilla learner. The metalearner is required to generalize out-of-distribution from these study examples to a test query (blue), which is analogous to the test set in the vanilla learner. By (meta)training on many such o.o.d. generalization problems (red, numbered 1, 2, 3), the metalearner can come to implement in its forward dynamics a separate learning algorithm with an inductive bias for compositionality. When a new o.o.d. generalization problem is sampled from the same distribution of tasks for (meta)testing (purple), the metalearner can generalize because it is familiar with such tasks (i.e., the task is i.i.d. with respect to its training data), even though the problem itself involves o.o.d. generalization.
@@
In this setting, we can distinguish the outer-loop (“base”) learning algorithm that determines the updates to the network’s weights during metalearning, and the inner-loop (“mesa-optimization”; Hubinger et al.,, 2021) or in-context learning algorithm that is implicitly implemented in the network’s forward dynamics (see Figure 5). The inner-loop algorithm does not require updates to the weights of the network to learn (i.e., it can still learn a new task when the weights are frozen).
Crucially, the inner-loop learning algorithm implemented in the model’s forward pass can have different properties than the outer-loop learning algorithm used to train the weights of the model (see Figure 5). For example, the inner-loop learning algorithm may have a completely different learning rate or updating procedure, or embody inductive biases that are not present in the outer-loop algorithm.
@@
This suggests a novel perspective on how a neural network might instantiate compositionality without incorporating explicit classical symbols: by (meta)learning from data to implement in its activation dynamics a learning algorithm with an inductive bias for compositionality. Even if the outer-loop learning algorithm used by neural networks lacks compositionality, as theorists such as Fodor and Pylyshyn, (1988) and Marcus, (1998) suggest, it is still possible that the inner-loop learning algorithm that a network comes to implement in its activation dynamics will exhibit the property of compositionality.
@@
Lake and Baroni, (2023), following up on earlier work (Lake,, 2019), applied this approach to compositional generalization problems such as the SCAN task (Lake and Baroni,, 2018), and showed that a metalearning transformer was capable of near-perfect accuracy on the task. The transformer was trained on a distribution of tasks, each of which was itself a compositional generalization problem. In particular, each task had a few samples from the grammar that were in principle sufficient to infer the rules that could be used to produce the correct answer on the test sample. The model was trained on some generalization tasks, and tested on ones that it had not seen. The results showed that the transformer was capable of metalearning to implement a compositional learning algorithm in its activation dynamics, achieving near-perfect accuracy on held-out compositional generalization tasks that were not seen during metalearning.
@@
These results demonstrate a novel strategy for endowing neural network models with the property of compositionality, and for overcoming the problem of o.o.d. generalization. After metalearning was complete, the model was given an o.o.d. problem in context, and its inner-loop learning algorithm exhibited an ability to generalize. However, from the perspective of the outer-loop learning algorithm, the problem was familiar (i.i.d.), because it was trained on similar o.o.d. generalization tasks during metalearning (see Figure 6). Thus, metalearning creates a scenario where two kinds of generalization can coexist simultaneously: the outer-loop learning algorithm is simply generalizing i.i.d., while the inner-loop learning algorithm is capable of generalizing o.o.d. because of the inductive biases imparted to it by outer-loop learning (see Figure 7).
Figure 7: Outer-loop generalization can be i.i.d., even if each generalization problem solved in the inner-loop is o.o.d. A metalearner can be trained on a distribution of tasks, where each task is itself a generalization problem like the SCAN task (see Figure 3). In this case, each sample (“episode”) in the outer loop corresponds to an o.o.d. generalization problem that must be solved by the inner-loop learning algorithm. The metalearner can be tested on new tasks from the same distribution (i.e., the outer-loop generalization problem is i.i.d.), where each new task is itself an o.o.d. generalization problem. In this figure, the colors are coordinated with Figure 6, where red and purple represent train and test samples in the outer loop, while green and blue represent the train and test samples in the inner loop.
@@
Lake and Baroni, (2023) conclude that although their metalearner could generalize on o.o.d. problems that were similar to those seen during metalearning, it could not generalize to completely new kinds of problems (i.e., ones that were o.o.d. with respect to the metalearning dataset). One way of interpreting this finding is to maintain that the kind of compositionality that humans exhibit is truly o.o.d., so the model’s failure shows that it is inadequate as a model of human compositionality. However, another interpretation invites a reexamination of human compositionality and suggests that it too could be a property of an inner-loop learning algorithm, in this case implemented by the human brain (Russin et al., 2024a, ). This would challenge the claim that humans are capable of generalizing compositionally in completely novel domains, instead establishing the weaker claim that humans generalize o.o.d., but only in specific domains determined either by evolutionary history or by prior learning experiences.
4.4 In-Context Compositionality via Pretraining

### v2

@@ 3.1 Operationalizing Compositionality
However, recent work has investigated methods that endow neural networks with inductive biases that would facilitate their adaptation to these shifts and allow them to replicate the behavioral signatures of compositionality. In the following, we’ll review recent work suggesting that neural networks can acquire such inductive biases via metalearning, or learning to learn. In the metalearning setting (Lake and Baroni,, 2023), an inductive bias is imparted to a neural network by training it on a specific distribution of tasks (e.g., compositional generalization tasks), endowing it with an ability to learn new tasks in specific ways. Similarly, when neural networks such as LLMs are pretrained on a large amount of data, they can acquire an emergent ability to learn new tasks given in context (Brown et al.,, 2020; Bubeck et al.,, 2023).
@@
In the following, we will review key findings from recent studies exploring compositionality in each of these settings, highlighting how they challenge the traditional view of compositionality in neural networks. Our review will largely focus on language, as this has been the domain in which compositionality has traditionally played a central explanatory role (see section 2.1). However, there has also been a recent surge of interest in compositionality in vision (Lepori et al.,, 2023; Zhou et al.,, 2024), multi-modal models (Castro et al.,, 2024; Hsieh et al.,, 2023; Lewis et al.,, 2023; Ma et al., 2023b, ; Ma et al., 2023a, ; Mitra et al.,, 2023; Ossowski et al.,, 2024; Shukor et al.,, 2023), and in reinforcement learning (Bakirtzis et al.,, 2022; Jothimurugan et al.,, 2021; Liu and Frank,, 2022; Žikelić et al.,, 2023), where similar issues have been explored. Furthermore, while we have chosen to focus our review on metalearning and LLMs, it should be noted that much recent work has pursued other approaches. For example, researchers have investigated whether certain architectural inductive biases can facilitate compositional generalization (Altabaa et al.,, 2023; Csordás et al.,, 2022; Kriete et al.,, 2013; Palangi et al.,, 2017; Rougier et al.,, 2005; Russin et al.,, 2020; Soulos et al.,, 2023; Schlag et al.,, 2019; Webb et al.,, 2024, 2023), explored techniques involving data augmentation (Andreas,, 2020; Cazzaro et al.,, 2024; Chai et al.,, 2023; Jiang et al.,, 2022) and the injection of structural knowledge via additional supervision or auxiliary tasks (Chakravarthy et al.,, 2022; Jiang and Bansal,, 2021), and investigated the limits of standard transformers on compositional tasks (Dziri et al.,, 2023; Zhou et al., 2023b, ).
@@
After the review, we discuss how these new approaches may inform our understanding of human compositionality, including their implications for its neural mechanisms and development. We conclude with a discussion of the ‘mere implementation’ objection, the second horn of Fodor and Pylyshyn’s dilemma.
3.2 Metalearning
In metalearning (Binz et al.,, 2023; Griffiths et al.,, 2019; Sandbrink and Summerfield,, 2024; Wang et al.,, 2018; Wang,, 2020), a model is trained not on a single task but on a distribution of tasks, giving it the opportunity to learn how to learn new tasks more rapidly and to leverage regularities among tasks to generalize in specific ways. Metalearning can impart specific inductive biases to a model. For example, a metalearning agent trained on a distribution of navigation tasks may learn exploration strategies that are specifically effective for that distribution, thus allowing it to learn more efficiently how to navigate in similar environments.
Importantly, these inductive biases are imparted to the model without specifying any particular algorithmic features in advance (McCoy et al.,, 2020). Instead, the researcher can simply define the set of learning problems at which the model should excel and rely on the optimization process during metalearning to discover whatever inductive biases are useful for accomplishing these behaviors.
@@ 3.2 Metalearning
A number of metalearning methods have been developed for use with neural networks (Bengio et al.,, 1991, 1995; Duan et al.,, 2016; Finn et al.,, 2017; Hospedales et al.,, 2020; Schmidhuber,, 1987). In many of these methods, the model learns to implement a separate “inner-loop” or “in-context” learning (ICL) algorithm in its forward activation dynamics, distinct from the “outer-loop” or “in-weight” learning (IWL) algorithm used to train the network in the first place. In this case, the network’s in-context learning algorithm “learns” within the flow of information from its inputs to its outputs (Duan et al.,, 2016; Lake,, 2019; Lake and Baroni,, 2023; Santoro et al.,, 2016; Wang et al.,, 2018). The network’s weights are trained such that the forward pass of the model can take some study examples as inputs in context and learn from them to produce the correct generalization behavior on a test query. For example, a sequence-processing network such as an RNN or a transformer can be trained on inputs that consist of (1) a set of study examples along with their labels, and (2) a single test query whose label is not given, but must be inferred. In this case, the network is trained to produce a prediction about the label corresponding to the test query, conditioned on the set of training samples and their labels. The model is then evaluated by freezing its weights and testing on new study examples and a new test query drawn from another task.
@@
In this setting, we can distinguish the IWL algorithm that determines the updates to the network’s weights during metalearning, and the ICL algorithm that is implicitly implemented in the network’s activation dynamics. The ICL algorithm does not require updates to the weights of the network to learn (i.e., it can still learn a new task when the weights are frozen). Crucially, these two distinct learning algorithms can have different properties. For example, the ICL algorithm may have a completely different updating procedure or effective learning rate, or may embody inductive biases that are not present in the original IWL algorithm.
@@
This suggests a novel perspective on how a neural network might instantiate compositionality without incorporating explicit classical symbols: by (meta)learning an ICL algorithm with an inductive bias for compositional generalization. Even if the usual IWL algorithm used by neural networks lacks compositionality, as theorists such as Fodor and Pylyshyn, (1988) and Marcus, (1998) suggest, it is still possible that the ICL algorithm that a network comes to implement in its activation dynamics will exhibit compositionality.
@@
Lake and Baroni, (2023), following up on earlier work (Lake,, 2019), applied this approach to compositional generalization problems such as the SCAN task (Lake and Baroni,, 2018), and showed that a metalearning transformer was capable of near-perfect accuracy on the task. The transformer was trained on a distribution of tasks, each a compositional generalization problem in its own right. In particular, each task had a few samples from the grammar that were in principle sufficient to infer the rules and produce the correct answer on the test sample. The model was trained on some generalization tasks, and tested on ones that it had not seen. The results showed that the transformer was capable of metalearning to implement a compositional learning algorithm in its activation dynamics, achieving near-perfect accuracy on held-out compositional generalization tasks that were not seen during metalearning.
@@
These results demonstrate a novel strategy for endowing neural network models with the property of compositionality, and for overcoming the problem of o.o.d. generalization. After metalearning was complete, the model was given an o.o.d. problem in context, and its ICL (inner-loop) algorithm exhibited an ability to generalize. However, from the perspective of the IWL (outer-loop) algorithm, the problem was familiar (i.i.d.), because it was trained on similar o.o.d. generalization tasks during metalearning. Thus, metalearning creates a scenario where two kinds of generalization can coexist simultaneously: the IWL algorithm is simply generalizing i.i.d., while the ICL algorithm is capable of generalizing o.o.d. This o.o.d. generalization is possible because of the inductive biases imparted to the ICL algorithm throughout metalearning.
@@
Lake and Baroni, (2023) conclude that although their metalearner could generalize on o.o.d. problems that were similar to those seen during metalearning, it could not generalize to completely new kinds of problems (i.e., ones that were o.o.d. with respect to the metalearning dataset). One way of interpreting this finding is to maintain that the kind of compositionality that humans exhibit is truly o.o.d., so the model’s failure shows that it is inadequate as a model of human compositionality. However, another interpretation invites a reexamination of human compositionality and suggests that it too could be a property of an ICL algorithm, in this case implemented by the human brain (Russin et al., 2024a, ). This would challenge the claim that humans are capable of generalizing compositionally in completely novel domains, instead establishing the weaker claim that humans generalize o.o.d., but only in specific domains determined either by evolutionary history or by prior learning experiences.
3.3 Large Language Models